# Time-dependent flux backgrounds in type IIB supergravity

### v1

@@ 7. Completing the tabulation of solvable subcases
• Now we discuss flux combinations within this specific ansatz for which the supergravity equations become an overdetermined system:
@@ 8. Miscellaneous remarks
In the time-dependent backgrounds in (38) constructed with nontrivial self-dual $\tilde{F}_{5}$ featuring $[3,2]$ and $[1,4]$ components and vanishing $G_{3}$, the axiodilaton remains an unfixed constant. [^48] The counting in the preceding discussion shows that this is the only case where $\tau$ can be kept constant. Moreover, the supergravity equations do not fix its value. Therefore, in this backgrounds the string coupling ($\textrm{Im}\tau=g_{s}^{-1}$) can be tuned to arbitrarily small values parametrically. This may be of particular importance for a worldsheet description of such backgrounds, as recently demonstrated by Cho:2023mhw for warped toroidal flux compactification for the type IIB flux configurations that allow arbitrarily weak string coupling.
@@ Acknowledgements.
ARK is supported by the Czech Science Foundation GAČR grant "Dualities and higher derivatives" (GA23-06498S) and thanks the CERN theory group and Irene Valenzuela for the hospitality at CERN. RM is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12247103.
@@ D.1 $G_{3}=0,\ \textrm{self-dual}\ \tilde{F}_{5},\ \tau=\tau_{0}$
The scale factor exponents are given in (38). Note that, due to the relation among the integration constants, no pair among $c_{3},c_{5},c_{7}$ can simultaneously vanish since $c_{2}^{2}>0$. For an admissible choice of $\beta$ (i.e., $0<e^{8\beta}<8c_{2}^{2}$) such that $e^{8\beta}$ is monotonically decreasing in time, we need to choose $c_{3},c_{5},c_{7}<0$ to obtain growing scale factors from nonzero initial values. [^50] Note that $\coth^{-1}[z]>0$ and monotonically decreasing with $z$ for $z>1$. See also footnote 16. To ensure a nonzero initial value of $e^{6\beta+\sigma}$, a nonvanishing initial $\dot{\beta}$ is also required.

### v2

@@ 7. Completing the tabulation of solvable subcases
• Now we discuss flux combinations within this specific ansatz (1) for which the supergravity equations become an overdetermined system:
@@ 8. Miscellaneous remarks
In the time-dependent backgrounds in (38) constructed with nontrivial self-dual $\tilde{F}_{5}$ featuring $[3,2]$ and $[1,4]$ components and vanishing $G_{3}$, the axiodilaton remains an unfixed constant. [^51] The counting in the preceding discussion shows that this is the only case where $\tau$ can be kept constant. Moreover, the supergravity equations do not fix its value. This is the only case in which $\textrm{Im}\tau$ cannot be traded with $\beta(t)$ as the undetermined function. Therefore, in this backgrounds the string coupling ($\textrm{Im}\tau=g_{s}^{-1}$) can be tuned to arbitrarily small values parametrically. This may be of particular importance for a worldsheet description of such backgrounds, as recently demonstrated by Cho:2023mhw for warped toroidal flux compactification for the type IIB flux configurations that allow arbitrarily weak string coupling.
@@
Lastly, we would like to point out that it would be interesting to allow for off-diagonal components (i.e., $\mu m$ terms) in the 10D metric (1) in order to obtain a more general class of type IIB backgrounds. As long as the $x$- and $y$-dependences of such components factorize, integrating over the $y$-coordinates and studying the resulting 4D effective theory should not involve further conceptual subtleties. Moreover, if such components depend only on $t$, the 4D spacetime component may still admit FLRW isometries in some cases. The analysis of section 4 could also be extended to incorporate such off-diagonal terms in (73).
@@ Acknowledgements.
We thank Andrew R. Frey for reading an earlier version of the manuscript and for providing valuable comments. We also thank Evan McDonough and Keshav Dasgupta for interesting comments and correspondence on the earlier version. ARK is supported by the Czech Science Foundation GAČR grant "Dualities and higher derivatives" (GA23-06498S) and thanks the CERN theory group and Irene Valenzuela for the hospitality at CERN. RM is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12247103.
@@ D.1 $G_{3}=0,\ \textrm{self-dual}\ \tilde{F}_{5},\ \tau=\tau_{0}$
The scale factor exponents are given in (38). Note that, due to the relation among the integration constants, no pair among $c_{3},c_{5},c_{7}$ can simultaneously vanish since $c_{2}^{2}>0$. For an admissible choice of $\beta$ (i.e., $0<e^{8\beta}<8c_{2}^{2}$) such that $e^{8\beta}$ is monotonically decreasing in time, we need to choose $c_{3},c_{5},c_{7}<0$ to obtain growing scale factors from nonzero initial values. [^53] Note that $\coth^{-1}[z]>0$ and monotonically decreasing with $z$ for $z>1$. See also footnote 17. To ensure a nonzero initial value of $e^{6\beta+\sigma}$, a nonvanishing initial $\dot{\beta}$ is also required.